# Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs): learning neural networks for designing neutral inclusions

### v1

@@ Abstract
We focus on designing and solving the neutral inclusion problem via neural networks. The neutral inclusion problem has a long history in the theory of composite materials, and it is exceedingly challenging to identify the precise condition that precipitates a general-shaped inclusion into a neutral inclusion. Physics-informed neural networks (PINNs) have recently become a highly successful approach to addressing both forward and inverse problems associated with partial differential equations. We found that traditional PINNs perform inadequately when applied to the inverse problem of designing neutral inclusions with arbitrary shapes. In this study, we introduce a novel approach, Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs), which integrates complex analysis techniques into PINNs. This method exhibits strong performance in solving forward-inverse problems to construct neutral inclusions of arbitrary shapes in two dimensions, where the imperfect interface condition on the inclusion’s boundary is modeled by training neural networks. Notably, we mathematically prove that training with a single linear field is sufficient to achieve neutrality for untrained linear fields in arbitrary directions, given a minor assumption. We demonstrate that CoCo-PINNs offer enhanced performances in terms of credibility, consistency, and stability.
@@ 1 Introduction
Physics-informed neural networks (PINNs)(Raissi et al., 2019; Karniadakis et al., 2021) are specialized neural networks designed to solve partial differential equations (PDEs). Since their introduction, PINNs have been successfully applied to a wide range of PDE-related problems (Cuomo et al., 2022; Hao et al., 2023; Wu et al., 2024). A significant advantage of using PINNs is their versatile applicability to different types of PDEs and their ability to deal with PDE parameters or initial/boundary constraints while solving forward problems (Akhound-Sadegh et al., 2023; Cho et al., 2023; Rathore et al., 2024; Lau et al., 2024). The conventional approach to solving inverse problems with PINNs involves designing neural networks that converge to the parameters or constraints to be reconstructed, which are typically modeled as constants or functions. We refer to this methodology as classical PINNs. Numerous successful outcomes in solving inverse problems using PINNs have been reported. See, for example, Chen et al. (2020); Jagtap et al. (2022); Haghighat et al. (2021). However, as the complexity of the PDE-based inverse problem increases, the neural networks may require additional design to represent the parameters or constraints accurately. For instance, Pokkunuru et al. (2023) utilized Bayesian approach to design the loss function, Guo et al. (2022) used Monte Carlo approximation to compute the fractional derivatives, Xu et al. (2023) adopted multi-task learning method to weight losses and also presented the forward-inverse problem combined neural networks, and Yuan et al. (2022) propose the auxiliary-PINNs to solve the forward and inverse problems of integro-differential equations. This increase in network complexity can significantly escalate computational difficulties and the volume of data necessary for training PINNs. Moreover, the direct approach to approximate the reconstructing parameters by neural networks enables too much fluent representation ability. This alludes to the fact that the conventional approach is inadequate depending on the problems due to the intrinsic ill-posedness of inverse problems.
@@
Figure 1.1: Inclusion makes perturbation.
@@
Inclusions with different material features from the background medium commonly cause perturbations in applied fields when they are inserted into the medium. Problems analyzing and manipulating the effects of inclusions have gained significant attention due to their fundamental importance in the modeling of composite materials, particularly in light of rapid advancements and diverse applications of these materials. Specific inclusions, referred to as neutral inclusions, do not disturb linear fields (see Figure 1.2). The neutral inclusion problem has a long and established history in the theory of composite materials (Milton, 2002). Some of the most well-known examples include coated disks and spheres (Hashin, 1962; Hashin & Shtrikman, 1962), as well as coated ellipses and ellipsoids (Grabovsky & Kohn, 1995; Kerker, 1975; Sihvola, 1997). The primary motivation for studying neutral inclusions is to design reinforced or embedded composite materials in such a way that the stress field remains unchanged from that of the material without inclusions and avoids stress concentration. Extensive research has been conducted on neutral inclusions and related concepts, such as invisibility cloaking involving wave propagation, in fields including acoustics, elasticity, electromagnetic waves within the microwave range Alù & Engheta (2005); Ammari et al. (2013); Landy & Smith (2013); Liu et al. (2017); Zhou & Hu (2006, 2007); Zhou et al. (2008); Yuste et al. (2018).
Figure 1.2: Neutral inclusions with circular shapes.
@@
Designing neutral inclusions with general shapes presents an inherent challenge. In the context of the conductivity problem, which is the focus of this paper, mathematical theory shows that only coated ellipses and ellipsoids can maintain neutral properties for linear fields in all directions (Kang & Lee, 2014; Kang et al., 2016; Milton & Serkov, 2001). In contrast, non-elliptical shapes can remain neutral for just a single linear field direction (Jarczyk & Mityushev, 2012; Milton & Serkov, 2001). To address the difficulty of designing general shaped neutral inclusions, relaxed versions of the problem have been studied (Choi et al., 2023; Kang et al., 2022; Lim & Milton, 2020). Differently from the above examples, where perfectly bonding boundaries are assumed, imperfect interfaces introduce discontinuities in either the flux or potential boundary conditions in PDEs. Ru (1998) found interface parameters for typical inclusion shapes in two-dimensional elasticity for typical inclusion shapes. Benveniste & Miloh (1999) found neutral inclusions under a single linear field. The interface parameters, which characterize these discontinuities, may be non-constant functions defined along the boundaries of the inclusions so that, theoretically, the degree of freedom of the interface parameters is infinite. Hence, we expect to overcome the inherent challenge of designing neutral inclusions with general shapes by considering neutral inclusions with imperfect interfaces. A powerful technique for dealing with planar inclusion problems of general shapes has been to use conformal mappings and to define orthogonal curvilinear coordinates (Movchan & Serkov, 1997; Cherkaev et al., 2022; Ammari & Kang, 2004; Jung & Lim, 2021), where the existence of the conformal mapping for a simply connected bounded domain is mathematically guaranteed by the Riemann mapping theory. Using these coordinates, Kang & Li (2019); Choi & Lim (2024) constructed weakly neutral inclusions that yield zero coefficients for leading-order terms in PDEs solution expansions (also refer to Milton (2002); Choi et al. (2023); Lim & Milton (2020) for neutral inclusion problems using the conformal mapping technique). However, such asymptotic approaches cannot achieve complete neutrality within this framework. Moreover, the requirement for analytic asymptotic expressions poses limitations on the generalizability of this approach.
@@
In this paper, by adopting a deep learning approach approach, we focus on precise values of the solution rather than asymptotic ones. Unlike the asymptotic approaches in Kang & Li (2019); Choi & Lim (2024), the proposed method does not rely on an analytical expansion formula and incorporates the actual solution values directly into the loss function design. More precisely, we introduce a novel forward-inverse PINN framework by combining complex analysis techniques into PINNs, namely Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs). We define the loss function to include the evaluations of the solutions at sample points exterior of inclusions. Furthermore, we leverage the conformal mapping to effectively sample collocation points for PDEs involving general shaped inclusions. We found that classical PINNs–which treat the interface parameters as functions approximated by neural networks–perform inadequately when applied to designing imperfect parameters for achieving neutrality. Instead, we propose training the Fourier series coefficients of the imperfect parameters, rather than approximating the function. We test the performance of our proposed method in finding the forward solution by using the analytical mathematical results for the forward solution presented in Choi & Lim (2024), where theoretical direct solutions are expressed as products of infinite dimensional matrices whose entries depends on the expansion coefficients of interface parameter. Additionally, we leverage these analytical results to explain that why it is possible to train the PINN for the neutral inclusion using only a single applied field (see Subsection 3.2).
@@
It is noteworthy that by utilizing Fourier series expansions to represent the inverse parameters, CoCo-PINNs offer deeper analytic insights into these parameters, making the solutions not only more accurate but also more explainable. Furthermore, our method requires no training data for the neutral inclusion problem due to its unique structure, where constraints at exterior points effectively serve as data. An additional remarkable feature is that our proposed method has proven effective in identifying optimal inverse parameters that are valid for general first-order background fields including those not previously trained. This impressive result is supported by a rigorous mathematical analysis.
@@
  • We developed a novel approach to PINNs, namely CoCo-PINNs, which have been shown to offer enhanced credibility, consistency, and stability compared to classical PINNs.
  • We have adopted the exterior conformal mapping in the PINNs to make it train the problem corresponding to the arbitrarily shaped domains.
  • Due to the nature of the neutral inclusion problem, it can be mathematically shown that once training with a given background solution $H(x)=x_{1}$ or $x_{2}$, a similar effect can be obtained for arbitrary linear background solutions $H(x)=ax_{1}+bx_{2}$ for any $a,b\in\mathbb{R}$ (see section 3.2 for more details).
@@ 2 Inverse problem of neutral inclusions with imperfect conditions
We set $x=(x_{1},x_{2})$ to be a two-dimensional vector in $\mathbb{R}^{2}$. On occasion, we regard $\mathbb{R}^{2}\cong\mathbb{C}$, whereby $x=x_{1}+ix_{2}$ will be used. We assume that $\Omega\subsetneq\mathbb{C}$ is a nonempty simply connected bounded domain with an analytic boundary (refer Appendix B). We assume the interior region $\Omega$ has the constant conductivity $\sigma_{c}$ while the background medium $\mathbb{R}^{2}\setminus\overline{\Omega}$ has the constant conductivity $\sigma_{m}$. We set $\sigma=\sigma_{c}\chi_{\Omega}+\sigma_{m}\chi_{\mathbb{R}^{2}\setminus\overline{\Omega}},$ where $\chi$ is the characteristic function. We further assume that the boundary of $\partial\Omega$ is not perfectly bonding, resulting in a discontinuity in the potential. This discontinuity is represented by a nonnegative real-valued function $p$ on $\partial\Omega$, referred to as the interface parameter or interface function. Specifically, we consider the following potential problem:
$\begin{cases}\nabla\cdot\sigma\nabla u=0&\text{in }\mathbb{R}^{2},\\[2.84526pt] p(u|^{+}-u|^{-})=\sigma_{m}\frac{\partial u}{\partial\nu}\big{|}^{+}&\text{on }\partial\Omega,\\[2.84526pt] \sigma_{m}\frac{\partial u}{\partial\nu}\big{|}^{+}=\sigma_{c}\frac{\partial u}{\partial\nu}\big{|}^{-}&\text{on }\partial\Omega,\\[2.84526pt] (u-H)(x)=O(|x|^{-1})&\text{as }|x|\to\infty,\end{cases}$ (2.1)
@@
We define $\Omega$ as a neutral inclusion for the imperfect interface problem (2.1) if $(u-H)(x)=0$ for all $x$ in the exterior region $\mathbb{R}^{2}\setminus\overline{\Omega}$ where $H(x)$ is any arbitrary linear fields, i.e., $H(x)=ax_{1}+bx_{2}$ for any $a,b\in\mathbb{R}$.
We explore the development of neural networks to find a specific interface function $p$ that makes $\Omega$ a neutral inclusion, given the inclusion $\Omega$ along with the conductivities $\sigma_{c}$ and $\sigma_{m}$, while simultaneously providing the forward solution $u$.
@@ 2.1 Series solution for the governing equation via conformal mapping
By the Riemann mapping theorem (see Appendix B), there exists a unique $\gamma>0$ and conformal mapping $\Psi$ from $D=\{w\in\mathbb{C}:|w|>\gamma\}$ onto $\mathbb{C}\setminus\overline{\Omega}$ such that $\Psi(\infty)=\infty$, $\Psi^{\prime}(\infty)=1$, and
$\Psi(w)=w+a_{0}+\frac{a_{1}}{w}+\frac{a_{2}}{w^{2}}+\cdots.$ (2.2)
We set $\rho_{0}=\ln\gamma$. We use modified polar coordinates $(\rho,\theta)\in[\rho_{0},\infty)\times[0,2\pi)$ via $z=\Psi(w)=\Psi(e^{\rho+i\theta})$. One can numerically compute the conformal mapping coefficients $\gamma$ and $a_{n}$ for a given domain $\Omega$ (Jung & Lim, 2021; Wala & Klöckner, 2018). In the following, we assume $\Psi$ is given. Choi & Lim (2024) obtained the solution expression for the solution $u$ using the geometric function theory for complex analytic functions for a given arbitrary analytic domain $\Omega$:
Theorem 2.1 (Analytic solution formula).
Let $F_{m}(z)$ be the Faber polynomials associated with $\Omega$ and the applied field is given by $H(z)=\mbox{Re}\left(\sum_{m=1}^{\infty}\alpha_{m}F_{m}(z)\right)$, the solution $u$ satisfies
$(u-H)(z)=\mbox{Re}\left(\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}s_{mn}w^{-n}\right)\quad\text{for }|w|>\gamma,$ (2.3)
$\bm{s}:=\{s_{mn}\}_{m,n\geq 1}=-\bm{\alpha}\widetilde{A}_{1}-\overline{\bm{\alpha}}\widetilde{A}_{2},$ (2.4)
where $\bm{\alpha}$ is a diagonal matrix whose entries are $\alpha_{m}$, and $\mbox{Re}(\cdot)$ denotes the real part of a complex number. The matrices $\widetilde{A}_{1}$ and $\widetilde{A}_{2}$ are determined by the conductivities, the shape of inclusion, and the coefficients of the expansion formula of the interface function $p(z)$; refer to Appendix B for the concept of Faber polynomials and explicit definitions.
From Theorem 2.1, the exact solution $u$ to eq. 2.1 can be obtained analytically. We denote this solution as $u_{p}$ for comparison with the trained forward solution by neural networks. This analytic solution $u_{p}$ will be used to evaluate the credibility of the PINNs forward solver, as discussed in Section 4.2.
@@ 2.2 Series representation of the interface function
In this subsection, based on complex analysis, we present the series expansion formula for designing the CoCo-PINNs. We assume that the interface function $p(x)$ is nonnegative, bounded, and continuous on $\partial\Omega$. By introducing a parametrization $x(\theta)$ of $\partial\Omega$ with $\theta\in[0,2\pi)$, the interface function admits a Fourier series expansion with respect to $\theta$:
$p(x(\theta))=a_{0}+\textstyle\sum_{k\in\mathbb{N}}\left(a_{k}\cos(k\theta)+b_{k}\sin(k\theta)\right),\quad\theta\in[0,2\pi),$
where $a_{k}$ and $b_{k}$ are real constant coefficients. In particular, we take $x(\theta)=\Psi(w)$, $w=\gamma e^{i\theta}$, where $\Psi$ is the conformal mapping in eq. 2.2. In this case, we have
$p(w):=p(x(\theta))$ $=\mbox{Re}\left(p_{0}+p_{1}w+p_{2}w^{2}+p_{3}w^{3}+\cdots\right)$ (2.5)
$=p_{0}+p_{1}w+\overline{p_{1}}\gamma^{2}w^{-1}+p_{2}w^{2}+\overline{p_{2}}\gamma^{4}w^{-2}+\cdots,\quad|w|=\gamma,$
for some complex-valued constants $p_{k}$. Note that we can similarly express the boundary conditions in eq. 2.1 in terms of the variable $w$, enabling us to effectively address these boundary conditions.
@@
$p^{(n)}(w)=\mbox{Re}\left(\textstyle\sum_{k=0}^{n}p_{k}w^{k}\right).$ (2.6)
At this state, the reconstruction parameters are deduced to $p_{0},\cdots,p_{n}$, and this makes the inverse problem of determining the interface parameter over-determined by using the constraints $(u-H)(x)\approx 0$ for many sample points exterior of $\Omega$.

### v2

# Conformal mapping based Physics-informed neural networks for designing neutral inclusions

@@ Abstract
We address the neutral inclusion problem with imperfect boundary conditions, focusing on designing interface functions for inclusions of arbitrary shapes. Traditional Physics-Informed Neural Networks (PINNs) struggle with this inverse problem, leading to the development of Conformal Mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs), which integrate geometric function theory with PINNs. CoCo-PINNs effectively solve forward-inverse problems by modeling the interface function through neural network training, which yields a neutral inclusion effect. This approach enhances the performance of PINNs in terms of credibility, consistency, and stability.
keywords:
Physics-informed neural networks , Imperfect interface problem , Neutral inclusion , Geometric function theory
[…]
@@ 1 Introduction
Physics-informed neural networks (PINNs)[1, 2] are specialized neural networks designed to solve partial differential equations (PDEs). Since their introduction, PINNs have been successfully applied to a wide range of PDE-related problems [3, 4, 5]. A significant advantage of using PINNs is their versatile applicability to different types of PDEs and their ability to deal with PDE parameters or initial/boundary constraints while solving forward problems [6, 7, 8, 9]. Balancing the multiple loss terms in PINNs is crucial for efficient training. Recent works have proposed adaptive loss-weighting schemes for PINNs [52, 53], in connection with loss-scaling approaches for multitask networks such as GradNorm [51]. The conventional approach to solving inverse problems with PINNs involves designing neural networks that converge to the parameters or constraints to be reconstructed, which are typically modeled as constants or functions. We refer to this methodology as classical PINNs. Numerous successful outcomes in solving inverse problems using PINNs have been reported. See, for example, [10, 11, 12, 49, 50]. However, as the complexity of the PDE-based inverse problem increases, the neural networks may require additional design to represent the parameters or constraints accurately. For instance, [13] utilized a Bayesian approach to design the loss function, [14] used Monte Carlo approximation to compute the fractional derivatives, [15] adopted a multi-task learning method to weight losses and also presented the forward-inverse problem combined neural networks, and [16] proposed the auxiliary-PINNs to solve the forward and inverse problems of integro-differential equations. This increase in network complexity can significantly escalate computational difficulties and the volume of data necessary for training PINNs. Moreover, the direct approach to approximating the reconstructing parameters via neural networks admits an unconstrained function class, which may lead to instability. This alludes to the fact that the conventional approach is inadequate depending on the problems, due to the intrinsic ill-posedness of inverse problems.
@@
Figure 1: Inclusion-induced perturbation: The inclusion yields a perturbation in the background field, which depends on the interior conductivity, the shape of the inclusion, and the interface conditions.
@@
Inclusions with different material features from the background medium commonly cause perturbations in applied fields when they are inserted into the medium (see Figure 1). Problems analyzing and manipulating the effects of inclusions have gained significant attention due to their fundamental importance in the modeling of composite materials, particularly in light of rapid advancements and diverse applications of these materials. Specific inclusions, referred to as neutral inclusions, do not disturb linear fields (see Figure 2). The neutral inclusion problem has a long and established history in the theory of composite materials [17]. Some of the most well-known examples include coated disks and spheres [18, 19], as well as coated ellipses and ellipsoids [20, 21, 22]. The primary motivation for studying neutral inclusions is to design reinforced or embedded composite materials in such a way that the stress field remains unchanged from that of the material without inclusions and avoids stress concentration. Extensive research has been conducted on neutral inclusions and related concepts, such as invisibility cloaking involving wave propagation, in fields including acoustics, elasticity, electromagnetic waves within the microwave range [23, 24, 25, 26, 27, 28, 29, 30].
Figure 2: Neutral inclusions with circular shapes: We consider coated circular inclusions with three distinct conductivity regions under perfect bonding conditions, as well as single disk-shaped inclusions with imperfect interface conditions.
@@
Designing neutral inclusions with general shapes presents an inherent challenge. In the context of the conductivity problem, which is the focus of this paper, mathematical theory shows that only coated ellipses and ellipsoids can maintain neutral properties for linear fields in all directions [31, 32, 33]. In contrast, non-elliptical shapes can remain neutral for just a single linear field direction [34, 33]. To address the difficulty of designing general-shaped neutral inclusions, relaxed versions of the problem have been studied [35, 36, 37]. Different from the above examples, where perfectly bonding boundaries are assumed, imperfect interfaces introduce discontinuities in either the flux or potential boundary conditions in PDEs. [38] found interface parameters for typical inclusion shapes in two-dimensional elasticity for typical inclusion shapes. [39] found neutral inclusions under a single linear field. The interface parameters, which characterize these discontinuities, may be non-constant functions defined along the boundaries of the inclusions so that, theoretically, the degree of freedom of the interface parameters is infinite. Hence, we expect to overcome the inherent challenge of designing neutral inclusions with general shapes by considering neutral inclusions with imperfect interfaces. A powerful technique for dealing with planar inclusion problems of general shapes has been to use conformal mappings and to define orthogonal curvilinear coordinates [40, 41, 42, 43], where the existence of the conformal mapping for a simply connected bounded domain is mathematically guaranteed by the Riemann mapping theory. Using these coordinates, [44, 45] constructed weakly neutral inclusions that yield zero coefficients for leading-order terms in PDEs solution expansions (also refer to [17, 35, 37] for neutral inclusion problems using the conformal mapping technique). However, such asymptotic approaches cannot achieve complete neutrality within this framework. Moreover, the requirement for analytic asymptotic expressions poses limitations on the generalizability of this approach.
@@
In this paper, by adopting a deep learning approach, we focus on precise values of the solution rather than asymptotic ones. Unlike the asymptotic approaches in [44, 45], the proposed method does not rely on an analytical expansion formula and incorporates the actual solution values directly into the loss function design. More precisely, we introduce a novel forward-inverse PINNs framework by combining complex analysis techniques into PINNs, namely Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs). We define the loss function to include the evaluations of the solutions at sample points exterior to inclusions. Furthermore, we leverage the conformal mapping to effectively sample collocation points for PDEs involving general-shaped inclusions. We found that classical PINNs–which treat the interface parameters as functions approximated by neural networks–perform inadequately when applied to designing imperfect parameters for achieving neutrality. Instead, we propose training the Fourier series coefficients of the imperfect parameters, rather than approximating the function. We test the performance of our proposed method in finding the forward solution by using the analytical mathematical results for the forward solution presented in [45], where theoretical direct solutions are expressed as products of infinite-dimensional matrices whose entries depend on the expansion coefficients of the interface parameter.
@@
It is noteworthy that by utilizing Fourier series expansions to represent the inverse parameters, CoCo-PINNs offer deeper analytic insights into these parameters, making the solutions not only more accurate but also more explainable. Furthermore, our method requires no training data for the neutral inclusion problem due to its unique structure, where constraints at exterior points effectively serve as data. An additional remarkable feature is that our proposed method has proven effective in identifying optimal inverse parameters that are valid for general first-order background fields, including those not previously trained. This impressive result is supported by a rigorous mathematical analysis.
@@
  • 1.
    We developed a novel approach to PINNs, termed CoCo-PINNs, which demonstrates improved credibility, consistency, and stability compared to classical PINNs in solving forward–inverse problems for conductivity equations with imperfect interface conditions.
  • 2.
    We incorporated the exterior conformal mapping into the PINNs framework to enable training on problems defined over arbitrarily shaped domains.
  • 3.
    Due to the nature of the neutral inclusion problem and the linearity of the governing equations, training the model with only the two background solutions $H(x)=x_{1}$ and $H(x)=x_{2}$ ensures that the neutral effect extends to all linear background fields of the form $H(x)=ax_{1}+bx_{2}$, for any $a,b\in\mathbb{R}$ (see Section 3.2 for details).
@@ 2 Inverse problem of neutral inclusions with imperfect conditions
We set $x=(x_{1},x_{2})$ to be a two-dimensional vector in $\mathbb{R}^{2}$. On occasion, we regard $\mathbb{R}^{2}\cong\mathbb{C}$, whereby $x=x_{1}+ix_{2}$ will be used. We assume that $\Omega\subsetneq\mathbb{C}$ is a nonempty simply connected bounded domain with an analytic boundary (refer A). We assume the interior region $\Omega$ has the constant conductivity $\sigma_{c}$ while the background medium $\mathbb{R}^{2}\setminus\overline{\Omega}$ has the constant conductivity $\sigma_{m}$. We set $\sigma=\sigma_{c}\chi_{\Omega}+\sigma_{m}\chi_{\mathbb{R}^{2}\setminus\overline{\Omega}},$ where $\chi$ is the characteristic function. We further assume that the boundary of $\partial\Omega$ is not perfectly bonding, resulting in a discontinuity in the potential. This discontinuity is described by a nonnegative real-valued function $p(x)$ defined on $\partial\Omega$, referred to as the interface parameter or interface function. Specifically, we consider the following potential problem:
$\begin{cases}\nabla\cdot\sigma\nabla u=0&\text{in }\mathbb{R}^{2},\\[2.84526pt] p(u|^{+}-u|^{-})=\sigma_{m}\frac{\partial u}{\partial\nu}\big|^{+}&\text{on }\partial\Omega,\\[2.84526pt] \sigma_{m}\frac{\partial u}{\partial\nu}\big|^{+}=\sigma_{c}\frac{\partial u}{\partial\nu}\big|^{-}&\text{on }\partial\Omega,\\[2.84526pt] (u-H)(x)=O(|x|^{-1})&\text{as }|x|\to\infty,\end{cases}$ (1)
@@
We define $\Omega$ as a neutral inclusion for the imperfect interface problem 1 if (uH)(x)=0(u-H)(x)=0 for all $x$ in the exterior region $\mathbb{R}^{2}\setminus\overline{\Omega}$ where $H(x)$ is any arbitrary linear fields, i.e., $H(x)=ax_{1}+bx_{2}$ for any $a,b\in\mathbb{R}$.
We develop a neural network framework to identify an interface function $p$ that makes the inclusion $\Omega$ a neutral inclusion, given the geometry of $\Omega$ and the conductivities $\sigma_{c}$ and $\sigma_{m}$, while also computing the corresponding forward solution $u$.
@@ 2.1 Series solution for the governing equation via conformal mapping
By the Riemann mapping theorem (see A), there exists a unique $\gamma>0$ and conformal mapping $\Psi$ from $D=\{w\in\mathbb{C}:|w|>\gamma\}$ onto $\mathbb{C}\setminus\overline{\Omega}$ such that $\Psi(\infty)=\infty$, $\Psi^{\prime}(\infty)=1$, and
$\Psi(w)=w+a_{0}+\frac{a_{1}}{w}+\frac{a_{2}}{w^{2}}+\cdots.$ (2)
We set $\rho_{0}=\ln\gamma$. We use modified polar coordinates $(\rho,\theta)\in[\rho_{0},\infty)\times[0,2\pi)$ via $z=\Psi(w)=\Psi(\exp({\rho+i\theta}))$. One can numerically compute the conformal mapping coefficients $\gamma$ and $a_{n}$ for a given domain $\Omega$ [43, 46]. In the following, we assume $\Psi$ is given.
[45] derived an explicit analytic solution $u$ to 1, denoted by $u_{p}$, using geometric function theory for complex analytic functions, given an arbitrary analytic domain $\Omega$ and an interface function $p(x)$. This solution serves as the ground truth for evaluating the credibility of the neural network’s forward prediction $u_{\text{NN}}$, as discussed in Section 4.3.
@@ 2.2 Series representation of the interface function
In this subsection, we derive a series expansion formula for designing the CoCo-PINNs, based on complex analysis. We assume that the interface function $p(x)$ is nonnegative, bounded, and continuous on $\partial\Omega$. By introducing a parametrization $x(\theta)$ of $\partial\Omega$ for $\theta\in[0,2\pi)$, the interface function can be expanded into a Fourier series in terms of $\theta$:
$p(x(\theta))=\sum_{n=-\infty}^{\infty}p_{n}w^{n},\quad\text{with }|w|=\gamma,$
where $p_{n}$ are complex-valued coefficients. In particular, we set $x(\theta)=\Psi(w)$, where $w=\gamma\exp({i\theta})$ and $\Psi$ is the conformal mapping defined in 2. Since the interface function is real-valued, we have
$p_{-n}=\overline{p_{n}}\gamma^{2n}.$
We refer the reader to [45] for further details. In this case, we have
$p(w)$ $=p(x(\theta))=p_{0}+p_{1}w+\overline{p_{1}}\gamma^{2}w^{-1}+p_{2}w^{2}+\overline{p_{2}}\gamma^{4}w^{-2}+\cdots,$
for a real-valued constant $p_{0}$, some complex-valued constants $\{p_{k}\}_{k=1}^{\infty}$, and $|w|=\gamma$. Note that we can similarly express the boundary conditions in 1 in terms of the variable $w$, enabling us to effectively address these boundary conditions.
@@
$p^{(n)}(w)=\Re\left(\textstyle\sum_{k=0}^{n}p_{k}w^{k}\right).$
At this stage, the reconstruction parameters are reduced to $p_{0},\cdots,p_{n}$, and this makes the inverse problem of determining the interface parameter over-determined by using the constraints $(u-H)(x)\approx 0$ for many sample points exterior of $\Omega$.